A student borrows $1000 at 10% for six months to pay tuition. Find the total amount due in six months ?
Answer #1

PV =1000
Rate per six month =10%/2 =5%
Number of Periods =6/6 =1
FV=PV*(1+r)^n =1000*(1+10%/2)^1 =1050